Use the Gauss-Jordan Elimination method to solve the system of linear equations.

xa + 3xb + xc = 4

2xa + 2xb + xc = -1

2xa + 3xb + xc = 3



1
Expert's answer
2021-09-03T07:35:28-0400
"\begin{matrix}
 x_a+3x_b+x_c=4 \\
\\
 2x_a+2x_b+x_c=-1 \\ \\
 2x_a+3x_b+x_c=3 \\
\end{matrix}"

The augmented matrix is


"A=\begin{pmatrix}
 1 & 3 & 1 & & 4 \\
 2 & 2 & 1 & & -1 \\
 2 & 3 & 1 & & 3 \\
\end{pmatrix}"

"R_2=R_2-2R_1"


"\begin{pmatrix}
 1 & 3 & 1 & & 4 \\
 0 & -4 & -1 & & -9 \\
 2 & 3 & 1 & & 3 \\
\end{pmatrix}"

"R_3=R_3-2R_1"


"\begin{pmatrix}
 1 & 3 & 1 & & 4 \\
 0 & -4 & -1 & & -9 \\
 0 & -3 & -1 & & -5 \\
\end{pmatrix}"

"R_2=R_2\/(-4)"


"\begin{pmatrix}
 1 & 3 & 1 & & 4 \\
 0 & 1 & 1/4 & & 9/4 \\
 0 & -3 & -1 & & -5 \\
\end{pmatrix}"

"R_1=R_1-3R_2"


"\begin{pmatrix}
 1 & 0 & 1/4 & & -11/4 \\
 0 & 1 & 1/4 & & 9/4 \\
 0 & -3 & -1 & & -5 \\
\end{pmatrix}"

"R_3=R_3+3R_2"


"\begin{pmatrix}
 1 & 0 & 1/4 & & -11/4 \\
 0 & 1 & 1/4 & & 9/4 \\
 0 & 0 & -1/4 & & 7/4 \\
\end{pmatrix}"

"R_3=-4R_3"


"\begin{pmatrix}
 1 & 0 & 1/4 & & -11/4 \\
 0 & 1 & 1/4 & & 9/4 \\
 0 & 0 & 1 & & -7\\
\end{pmatrix}"

"R_1=R_1-R_3\/4"


"\begin{pmatrix}
 1 & 0 & 0 & & -1 \\
 0 & 1 & 1/4 & & 9/4 \\
 0 & 0 & 1 & & -7\\
\end{pmatrix}"

"R_2=R_2-R_3\/4"


"\begin{pmatrix}
 1 & 0 & 0 & & -1 \\
 0 & 1 & 0 & & 4 \\
 0 & 0 & 1 & & -7\\
\end{pmatrix}"

Then "x_a=-1, x_b=4, x_c=-7."


The solution is "(-1, 4, -7)."